# Learning Getting-Up Policies for Real-World Humanoid Robots

### v1

@@ III-C Stage I to Stage II Curriculum
The design of two-stage policy learning intrinsically constructs a hard-to-easy curriculum [7]. Stage I targets motion discovery in easier settings (weak regularization, no variations in terrain, same starting poses, simpler collision geometry). Once motions have been discovered, Stage II solves the task of making the learned motion deployable and generalizable. As our experiments will show, splitting the work into two phases is crucial for successful learning. Specifically, complexity increases from Stage I to Stage II in the following ways:
@@ III-C3 Control Regularization and Terrain Randomization
• Weak $\rightarrow$ strong control regularization. Weak control regularization in Stage I enables discovery of getting-up / rolling-over motion, while strong control regularization (via smoothness rewards, DoF velocity penalties, etc, see the full list in supplementary) in Stage II encourages more deployable action.
@@
TABLE I: Simulation results. We compare HumanUP with several baselines on the held-out split of our curated posture set $\mathcal{P}_{\text{supine}}$ and $\mathcal{P}_{\text{prone}}$ using full URDF. All methods are trained on the training split of our posture set $\mathcal{P}$, except for methods HumanUP w/o Stage II and w/o posture randomization. HumanUP solves task ❸ by solving task ❷ and task ❶ consecutively. We do not include the results of baseline 6 (HumanUP w/o Two-Stage Learning) as it cannot solve the task. Tao et al. [65] is trained to directly solving task ❸ as it does not have a rolling over policy. Sim2Real column indicates whether the method can transfer to the real world successfully. We tested all methods in the real world for which the Smoothness and Safety metrics are reasonable, and Sim2Real is false if deployment wasn’t successful. Metrics are introduced in Section V-C.
@@
HumanUP w/o Stage II 27.59 $\pm$ 0.28 1.23 $\pm$ 0.00 5.56 $\pm$ 0.36 0.45 $\pm$ 0.04 1213.07 $\pm$ 5.56 0.67 $\pm$ 4.71e-3 0.71 $\pm$ 2.17e-3
HumanUP w/o Full URDF 89.59 $\pm$ 0.29 0.82 $\pm$ 0.00 0.44 $\pm$ 0.01 0.08 $\pm$ 0.00 77.61 $\pm$ 0.86 0.92 $\pm$ 2.88e-5 0.75 $\pm$ 3.19e-5
@@ IV-B Simulation Configurations
We use Isaac Gym [52] for simulated training and evaluation. We use a URDF with simplified collision for stage-I training and the official whole-body URDF from Unitree [70] when training stage-II. To accurately model the numerous contacts between the humanoid and the ground, we use a high simulation frequency of 1000 Hz, while the low-level PD controller frequency operates at 50 Hz.
@@ V-B Baselines
    HumanUP w/ Hard Symmetry: Our policy trained using a humanoid with a symmetric controller. This symmetric controller follows the symmetry control principle of the G1 controller baseline described in real-world experiments, which leads to bilaterally symmetric control. We set all pitch DoFs actions to be the same between the left and the right DoFs, while flipping the directions of all the roll and yaw actions.
@@ V-C Metrics
• Task Success. We report task success rate Success (%) and Task Metrics: the head height (m) for the getting-up task, and cosine of the angle between the robot’s X-axis (sticking out to the front from the torso) and the gravity vector, for the rolling over task.
• Smoothness. We measure the Action Jitter (rad/$\text{s}^{3}$), DoF Pos Jitter (rad/$\text{s}^{3}$), and mean Energy ($N\cdot$ rad/s) for action smoothness evaluation. The jitter metrics are computed as the third derivative values [16], which indicates the coordination stability of body movements.
@@ V-D Results and Analysis
Table I presents evaluation results based on policies tested on held-out subsets of our curated initial posture set $\mathcal{P}$—10K samples each from $\mathcal{P}_{\text{supine}}$ and $\mathcal{P}_{\text{prone}}$. Fig. 4 shows the learning curve for the getting-up task, where the termination base height reflects the robot’s ability to lift its body, and body uprightness indicates whether it achieves a stable standing posture.
Figure 3: Real-world results. We evaluate HumanUP (ours) in several real-world setups that span diverse surface properties, including both man-made and natural surfaces, and cover a wide range of roughness (rough concrete to slippery snow), bumpiness (flat concrete to tiles), ground compliance (completely firm concrete to being swampy muddy grass), and slope (flat to about $10^{\circ}$). We compare HumanUP with G1’s built-in getting-up controller and our HumanUP w/o posture randomization (PR). HumanUP succeeds more consistently (78.3% vs 41.7%) and can solve terrains that the G1’s controller can’t.
@@ V-D1 Ignoring Torque / Control Limits Leads to Undeployable Policies
While [65] and HumanUP achieve similar success rates, the smoothness and safety metrics for [65] are significantly worse than HumanUP. For example, the average action jitter metric is nearly 18$\times$ higher than HumanUP. Actions from [65] are highly unstable and unsafe and thus cannot be safely deployed to the real robot. Furthermore, [65] learns a very fast getting-up motion that keeps jumping after getting up. See visualization [65]’s getting up motion in the supplementary material. A similar trend can be seen when comparing HumanUP to HumanUP w/o Stage II. While HumanUP w/o Stage II also solves the task to some extent, it achieves unsatisfying smoothness and safety metrics similar making it inappropriate for real-world deployment. Thus, the regularization imposed in Stage II is essential to making policies more amenable to Sim2Real transfer.
@@ V-D2 Importance of Learning via a Curriculum
Somewhat surprisingly, even though HumanUP w/o Full URDF was trained without the full URDF mesh, it generalizes fine when tested with the full URDF in simulation, as reported in Table I. However, we found poor transfer of this policy to the real world. It failed on all 5 trials on a simple flat terrain. We believe the poor real-world performance was because of the mismatch between the contact it was expecting and the contact that actually happened.
@@ V-D4 Posture randomization helps
HumanUP w/o posture randomization (PR) works much worse than HumanUP suggesting that PR is necessary for generalizable control.
@@
Compared to HumanUP w/ Hard Symmetry, HumanUP achieves better task success in Table I, particularly for the rolling-over task which is very difficult with symmetric commands.
@@
Figure 5: Getting up comparison with G1 controller. G1 controller uses a handcrafted motion trajectory, which can be divided into three phases, while our HumanUP learns a continuous and more efficient whole-body getting-up motion. Our HumanUP enables the humanoid to get up within 6 seconds, half of the G1 controller’s 11 seconds of control. (a), (b), and (c) record the corresponding mean motor temperature of the upper body, lower body, and waist, respectively. G1’s default controller’s execution causes the arm motors to heat up significantly, whereas our policy makes more use of the leg motors that are larger (higher torque limit of 83N as opposed to 25N for the arm motors) and thus able to take more load.
Figure 6: Qualitative examples of failure modes on grass slope and snow field. G1 controller isn’t able to squat on the sloping grass and slips on the slow. HumanUP policy is able to partially get up on both the slope and the snow but falls due to unstable feet placement on the slope and slippage on the snow.
@@ VI Real World Results
We compare our final HumanUP policy with 1) G1 controller and 2) a high-performing ablation of HumanUP (HumanUP w/o posture randomization). The G1 controller, that comes with the robot, tracks a hand-crafted trajectory in three-phase shown in Fig. 5: Phase 0 brings the robot to a canonical lying pose; Phase 1 first props up and then slides the torso forward using hands, followed by bending legs to squat; Phase 2 uses its waist to tilt up the torso to stand up from squatting. The motions in phase 1 and phase 2 are symmetric, and this default G1 controller only works for supine poses.
@@ VI-A Results
Fig. 3 presents the experimental results. Overall we find that HumanUP policies perform better than the G1 controller and the version of HumanUP without posture randomization (PR). We discuss the results and observed behavior further.
@@ VI-A1 Getting up from supine poses
The default G1 controller works under nominal conditions, i.e., firm, flat concrete ground with a reasonable friction value. However, it starts to fail on more challenging terrains. For the bumpier and rougher terrains (brick surface and stone tiles), the arms may get stuck between bumps causing failures. On slopes, the robot fails to squat or hoist itself up due to both the resistance of the grassland and the unstable squatting pose prone to falling caused by slopes. On the compliant ground, the robot gets destabilized due to the compliance of the ground. On slippery snow, the robot slips.
Both versions of HumanUP outperform the default G1 controller. Trained with terrain and domain randomization, they are robust to real-world variations such as slipperiness, bumps, and slopes. Dynamics randomization further enhances resilience to minor perturbations like slippage or ground compliance. The full method, incorporating posture randomization, performs better than the variant without it, as it is specifically trained to handle diverse initial configurations. Overall, HumanUP achieves a 78.3% success rate in getting up.
@@ VI-A2 Rolling over from prone to supine poses
Findings are similar for the rolling over task. As noted, the default controller can’t handle this situation. The full model exhibits more robust performance than the model trained without posture randomization. Rolling over seems to be easier than getting up, HumanUP achieves a 98.3% success rate for rolling over.
@@ VI-B Motion analysis
Fig. 5 show the motion and recorded motor temperatures for executions of the G1 controller and HumanUP policy.
@@ VI-B1 Motor temperature
Even when the G1 controller works, it uses the arms during Phase 1 of getting up. Fig. 5 (a) shows that the default controller’s execution causes the arm motors to heat up significantly more when compared to HumanUP execution. In contrast, our policy makes more use of the leg motors that are larger (higher torque limit of $83N$ as opposed to $25N$ for the arm motors) and thus able to take more load.
@@ VI-B2 Efficiency
In addition, we find that HumanUP gets the robot to stand successfully within about 6 seconds through a smooth and continuous motion, which is over 2$\times$ more efficient compared to the G1 controller which takes nearly 11 seconds.
@@ VI-C Failure Mode Analysis
Fig. 6 shows example failure modes for the G1 controller and HumanUP on challenging terrains. Fig. 6 (a) shows that the G1 controller tries to utilize the robot’s hands to squat, while the sloping ground prevents it from getting to the full squatting pose due to high friction and weak waist torques to move against the dumping tendency. In contrast, HumanUP manages to lift the body, while the sloping ground sometimes causes an unstable foot orientation. Fig. 6 (b) shows that on even more challenging terrains like snow fields, both G1 and HumanUP controllers may fail due to the slippery and deformable ground.
@@ VII Limitations
HumanUP has several limitations: 1) It relies on high-performance physics platforms like IsaacGym [52] to simulate contact-rich tasks such as getting up and rolling over. However, current robotics simulators lag behind those in animation and gaming, limiting the accuracy and efficiency of contact dynamics simulation, advancements like Genesis [6] and Mujoco Playground [75] are helpful. 2) The RL process in HumanUP is a under-specified problem [3, 13], making it difficult to ensure that the learned motion aligns precisely with human-like behavior. e.g., our discovered motion tends to raise hands for balance.
@@ VIII Discussion
In this paper, we tackle the problem of learning getting-up controllers for real-world humanoid robots. Different from locomotion tasks, getting up involves complex contact patterns that are not known apriori. We develop a two-stage solution for this problem based on reinforcement learning and sim-to-real. Stage I finds a solution under minimal constraints, while Stage II learns to track the trajectory discovered in Stage I under regularization on control and from varied starting poses and on varied terrains. We found this two stage strategy to be effective both in simulation and the real world. Specifically, it enabled us to get a real-world G1 humanoid to stand up from a supine pose and roll over from a supine pose to a prone pose on different terrains and from different starting poses. HumanUP achieves a higher success rate than the default G1 controller and expands the capabilities of the G1 robot.
We hope our learned policies for automatic fall recovery will be useful to researchers and practitioners, while our two stage learning framework may be helpful for other problems that require figuring out complex contact patterns.
Appendix A Material Statements
In our supplementary material, we provide a video demo presenting HumanUP, and this PDF for supplementary information including our reward designs, training details, and additional qualitative analysis.
Rewards
A-1 Rewards components in Stage I
Detailed reward components used in Stage I are summarized in Table II.
@@ A-1 Rewards components in Stage I
DoF position error $\mathds{1}(h_{\text{base}}\geq 0.8)\cdot\exp\left(-0.05\|{\boldsymbol{{d}}_{t}}-{\boldsymbol{{d}}_{t}}^{\text{default}}|\right)$0.75
@@
Feet height reward $\exp(-10\cdot\boldsymbol{h}^{\text{feet}})$ 2.5
Foot orientation $\sqrt{\lVert G_{xy}^{\text{feet}}\rVert}$0.5
@@
Base Gravity Exp $\frac{1}{2}\Big{(}\exp\big{(}-0.01(1-\cos\theta_{\text{left}})\big{)}+$ $\exp\big{(}-0.01(1-\cos\theta_{\text{right}})\big{)}\Big{)},$ $\cos\theta=\frac{\mathbf{g}^{\text{knee}}\cdot\mathbf{g}_{\text{target}}}{\|\mathbf{g}^{\text{base}}\|\|\mathbf{g}_{\text{base}}\|}$ 8
Knee Gravity Exp $\frac{1}{2}\Big{(}\exp\big{(}-0.01(1-\cos\theta_{\text{left}})\big{)}+$ $\exp\big{(}-0.01(1-\cos\theta_{\text{right}})\big{)}\Big{)},$ $\cos\theta=\frac{\mathbf{g}^{\text{knee}}\cdot\mathbf{g}_{\text{target}}}{\|\mathbf{g}^{\text{base}}\|\|\mathbf{g}_{\text{base}}\|}$ 8
Feet height reward $\exp(-10\cdot\boldsymbol{h}^{\text{feet}})$ 2.5
A-2 Rewards components in Stage II
Detailed reward components used in Stage II are summarized in Table III.

### v2

@@ III-C Stage I to Stage II Curriculum
The design of two-stage policy learning induces a hard-to-easy curriculum [7]. Stage I targets motion discovery in easier settings (weak regularization, no variations in terrain, same starting poses, simpler collision geometry). Once motions have been discovered, Stage II solves the task of making the learned motion deployable and generalizable. As our experiments will show, splitting the work into two phases is crucial for successful learning. Specifically, complexity increases from Stage I to Stage II in the following ways:
@@ III-C3 Control Regularization and Terrain Randomization
• Weak $\rightarrow$ strong control regularization. Weak control regularization in Stage I enables discovery of getting-up / rolling-over motion, while strong control regularization (via smoothness rewards, DoF velocity penalties, etc, see the full list in Sec. A.2) in Stage II encourages more deployable action.
@@
TABLE I: Simulation results. We compare HumanUP with several baselines on the held-out split of our curated posture set $\mathcal{P}_{\text{supine}}$ and $\mathcal{P}_{\text{prone}}$ using full URDF. All methods are trained on the training split of our posture set $\mathcal{P}$, except for methods HumanUP w/o Stage II and w/o posture randomization. HumanUP solves task ❸ by solving task ❷ and task ❶ consecutively. We do not include the results of baseline 6 (HumanUP w/o Two-Stage Learning) as it cannot solve the task. Tao et al. [70] is trained to directly solving task ❸ as it does not have a rolling over policy. Sim2Real column indicates whether the method can transfer to the real world successfully. We tested all methods in the real world for which the Smoothness and Safety metrics are reasonable, and Sim2Real is false if deployment wasn’t successful. Metrics are introduced in Sec. V-C.
@@
HumanUP w/o Stage II 27.59 $\pm$ 0.28 0.82 $\pm$ 0.00 5.56 $\pm$ 0.36 0.45 $\pm$ 0.04 1213.07 $\pm$ 5.56 0.67 $\pm$ 4.71e-3 0.71 $\pm$ 2.17e-3
HumanUP w/o Full URDF 89.59 $\pm$ 0.29 1.23 $\pm$ 0.00 0.44 $\pm$ 0.01 0.08 $\pm$ 0.00 77.61 $\pm$ 0.86 0.92 $\pm$ 2.88e-5 0.75 $\pm$ 3.19e-5
@@ IV-B Simulation Configurations
We use Isaac Gym [56] for simulated training and evaluation. We use a URDF with simplified collision for stage-I training and the official whole-body URDF from Unitree [75] for stage-II. To accurately model the numerous contacts between the humanoid and the ground, we use a high simulation frequency of 1000 Hz, while the low-level PD controller frequency operates at 50 Hz. More details can be found in Appendix C.
@@ V-B Baselines
    HumanUP w/ Hard Symmetry: Our policy trained using a humanoid with a symmetric controller. This symmetric controller follows the symmetry control principle of the manufacturer-provided controller baseline described in real-world experiments, which leads to bilaterally symmetric control. We set all pitch DoFs actions to be the same between the left and the right DoFs, while flipping the directions of all the roll and yaw actions.
@@ V-C Metrics
• Task Success. i) Task success rate Success (%): For the getting-up task, the robot’s head height must be $\geq$ 1.1m at termination, thus, the robot needs to continue to stand for success. For the rolling-over task, the cosine between the robot’s base, knee, and torso orientation and the target orientation when lying face up should be $\geq 0.9$, thus, the robot needs to move until lying face up. ii) Task Metrics: the head height (m) for the getting-up task, and the cosine of the angle between the robot’s torso X-axis (sticking out to the front from the torso) and the gravity, for the rolling-over task.
• Smoothness. We measure the Action Jitter ($\text{rad}/\text{s}^{3}$), DoF Pos Jitter ($\text{rad}/\text{s}^{3}$), and mean Energy ($N\cdot m\cdot\text{rad}/\text{s}$ ) for action smoothness evaluation [18]. The jitter metrics are computed as the third derivative values [16], which indicate the coordination stability of body movements.
@@ V-D Results and Analysis
Tab. I presents results based on policies tested on held-out subsets of our curated initial posture set $\mathcal{P}$, i.e. 10K val samples each from $\mathcal{P}_{\text{supine}}$ and $\mathcal{P}_{\text{prone}}$. Fig. 4 shows the learning curve for the getting-up task, where the termination base height reflects the robot’s ability to lift its body, and body uprightness indicates whether it achieves a stable standing posture.
Figure 3: Real-world results. We evaluate HumanUP (ours) in several real setups that span diverse surface properties, including both man-made and natural surfaces, and cover a wide range of roughness (rough concrete to slippery snow), bumpiness (flat concrete to tiles), ground compliance (completely firm concrete to being swampy muddy grass), and slope (flat to about $10^{\circ}$). We compare HumanUP with G1’s manufacturer-provided controller and HumanUP w/o posture randomization (PR). HumanUP succeeds more consistently (78.3% vs 41.7%) and can solve terrains that the manufacturer-provided controller can’t.
@@ V-D1 Ignoring Torque / Control Limits Leads to Undeployable Policies
While [70] and HumanUP achieve similar success rates, the smoothness and safety metrics for [70] are significantly worse than HumanUP. For example, the average action jitter metric is nearly 18$\times$ higher than HumanUP. Actions from [70] are highly unstable and unsafe and thus cannot be safely deployed to the real robot. Furthermore, [70] learns a very fast getting-up motion that keeps jumping after getting up. See visualization [70]’s getting up motion in Sec. B.1. A similar trend can be seen when comparing HumanUP to HumanUP w/o Stage II. While HumanUP w/o Stage II also solves the task to some extent, it achieves unsatisfying smoothness and safety metrics similar making it inappropriate for real-world deployment. Thus, the regularization imposed in Stage II is essential to making policies more amenable to Sim2Real transfer.
@@ V-D2 Importance of Learning via a Curriculum
Somewhat surprisingly, even though HumanUP w/o Full URDF was trained without the full URDF mesh, it generalizes fine when tested with the full URDF in simulation, as reported in Tab. I. However, we found poor transfer of this policy to the real world. It failed on all 5 trials on a simple flat terrain. We believe the poor real-world performance was because of the mismatch between the contact it was expecting and the contact that actually happened.
@@ V-D4 Posture randomization helps
HumanUP w/o posture randomization (PR) works much worse than HumanUP, suggesting that PR is necessary for generalizable control.
@@
Compared to HumanUP w/ Hard Symmetry, HumanUP achieves better task success in Tab. I, particularly for the rolling-over task, which is very difficult with symmetric commands.
@@
Figure 5: Getting up execution comparison with G1’s manufacturer-provided controller. The manufacturer-provided controller uses a handcrafted motion trajectory, which can be divided into three phases, while our HumanUP learns a continuous and more efficient whole-body getting-up motion. Our HumanUP enables the humanoid to get up within 6 seconds, half of the manufacturer-provided controller’s 11 seconds of control. (a), (b), and (c) record the corresponding mean motor temperature of the upper body, lower body, and waist, respectively. G1’s manufacturer-provided controller’s execution causes the arm motors to heat up significantly, whereas our policy makes more use of the leg motors that are stronger (higher torque limit of 83N as opposed to 25N for the arm motors) and thus able to take more load.
Figure 6: Qualitative examples of failure modes on grass slope and snow field. G1’s manufacturer-provided controller isn’t able to squat on the sloping grass and slips on the slope. HumanUP policy can partially get up on both the slope and the snow, but falls due to unstable foot placement on the slope and slippage on the snow.
@@ VI Real World Results
We compare our policy with 1) Manufacturer-provided Controller and 2) a high-performing ablation of HumanUP (HumanUP w/o posture randomization). The manufacturer-provided controller, which comes with the robot G1, tracks a hand-crafted trajectory in three phases shown in Fig. 5: Phase 0 brings the robot to a canonical lying pose; Phase 1 first props up and then slides the torso forward using hands, followed by bending legs to squat; Phase 2 uses its waist to tilt up the torso to stand up from squatting. Motions in phase 1 and phase 2 are symmetric, and this controller only works for supine poses.
@@ VI-A Results
Fig. 3 presents experimental results. Overall, we find that HumanUP policies perform better than the manufacturer-provided controller and HumanUP without posture randomization (PR). We discuss the results and observed behavior further.
@@ VI-A1 Getting up from supine poses
The manufacturer-provided controller works under nominal conditions, i.e., firm, flat concrete ground with a reasonable friction value. However, it starts to fail on more challenging terrains. For the bumpier and rougher terrains (brick surface and stone tiles), the arms may get stuck between bumps, causing failures. On slopes, the robot fails to squat or hoist itself up due to both the resistance of the grassland and the unstable squatting pose prone to falling caused by slopes. On the compliant ground, the robot gets destabilized. On slippery snow, the robot slips.
Both versions of HumanUP outperform the manufacturer-provided controller. Trained with terrain and domain randomization, they are robust to real-world variations such as slipperiness, bumps, and slopes. Dynamics randomization further enhances resilience to minor perturbations like slippage or ground compliance. The full method, incorporating posture randomization, performs better than the variant without it, as it is specifically trained to handle diverse initial configurations. Overall, HumanUP achieves a 78.3% getting-up success rate.
@@ VI-A2 Rolling over from prone to supine poses
Findings are similar for the rolling over task. As noted, the manufacturer-provided controller can’t handle this situation. The full model exhibits more robust performance than the model trained without posture randomization. Rolling over seems to be easier than getting up, HumanUP achieves a 98.3% success rate.
@@ VI-B Motion analysis
Fig. 5 shows the motion and motor temperatures for the manufacturer-provided controller and HumanUP policy.
@@ VI-B1 Motor temperature
The manufacturer-provided controller uses the arms during Phase 1 of getting up. Fig. 5 (a) shows that the default controller’s execution causes the arm motors to heat up significantly more when compared to HumanUP execution. Our policy makes more use of the leg motors that are stronger (higher torque limit of $83N$ as opposed to $25N$ for the arm motors) and thus able to take more load.
@@ VI-B2 Efficiency
HumanUP gets the robot to stand successfully within about 6 seconds through a smooth and continuous motion, which is over 2$\times$ more efficient than the manufacturer-provided controller, which takes nearly 11 seconds.
@@ VI-C Failure Mode Analysis
Fig. 6 shows example failure modes for the manufacturer-provided controller and HumanUP on challenging terrains. Fig. 6 (a) shows that the manufacturer-provided controller tries to utilize the robot’s hands to squat, while the sloping ground prevents it from getting to the full squatting pose due to high friction and weak waist torques to move against the dumping tendency. In contrast, HumanUP manages to lift the body, while the sloping ground sometimes causes an unstable foot orientation. Fig. 6 (b) shows that on even more challenging terrains like snow fields, both manufacturer-provided and HumanUP controllers may fail due to the slippery and deformable ground.
@@ VII Limitations
HumanUP has several limitations: 1) Motions discovered in Stage I could be incompatible with stronger control regularization used in Stage II. We didn’t encounter this issue in our experiments, possibly because of the weak control regularization used in Stage I and the use of 8×8$\times$8 × slower motion in Stage II. 2) HumanUP depends on high-performance physics simulators (IsaacGym [56]) running at high frequency (e.g., 1 kHz). Simulation speed and fidelity for more complex tasks involving perception and contacts remain a challenge. Recent advances such as Genesis [6], Mujoco Playground [80], and Roboverse [25] could help address these limitations. 3) The RL formulation in HumanUP is under-specified [3] and may lead to reward hacking [13], complicating precise alignment with natural human behaviors. For instance, our learned motions sometimes include unnatural hand raising for balance. 4) Extending HumanUP to handle more complex terrains like stairs or uneven surfaces remains under-explored, while humanoid robots may fall more easily on such terrains. Encouraging adaptive behaviors involving strong-arm usage on more powerful platforms may be useful to properly handle such situations.
@@ VIII Discussion
In this paper, we tackle the problem of learning getting-up controllers for real-world humanoid robots. Different from locomotion tasks, getting up involves complex contact patterns that are not known apriori. We develop a two-stage solution for this problem based on reinforcement learning and sim-to-real. Stage I finds a solution under minimal constraints, while Stage II learns to track the trajectory discovered in Stage I under regularization on control and from varied starting poses and on varied terrains. We found this two stage strategy to be effective both in simulation and the real world. Specifically, it enabled us to get a real-world G1 humanoid to stand up from a supine pose and roll over from a supine pose to a prone pose on different terrains and from different starting poses. HumanUP achieves a higher success rate than G1’s manufacturer-provided controller and expands the capabilities of the G1 robot.
We hope our learned policies for automatic fall recovery will be useful to researchers and practitioners, while our two-stage learning framework may be helpful for other problems that require figuring out complex contact patterns.
Acknowledgments
This material is based upon work supported by an NSF CAREER Award (IIS2143873) and a DURIP grant (N00014-23-1-2166). We also thank the Coordinated Science Laboratory for providing experimental space.
Appendix
[…]
1
@@ Acknowledgments
DoF position error $\mathds{1}(h_{\text{base}}\geq 0.8)\cdot\exp\left(-0.05\|{\boldsymbol{{d}}_{t}}-{\boldsymbol{{d}}_{t}}^{\text{default}}\|\right)$0.75
@@
Feet orientation $\sqrt{\lVert\mathbf{g}_{xy}^{\text{feet}}\rVert}$0.5
Feet height reward $\exp(-10\cdot\boldsymbol{h}^{\text{feet}})$ 2.5
@@
Base Gravity Error $1-\cos\theta_{\text{base}}$ $\cos\theta_{\text{base}}=\frac{\mathbf{g}^{\text{base}}\cdot\mathbf{g}^{\text{target}}}{\|\mathbf{g}^{\text{base}}\|\|\mathbf{g}^{\text{target}}\|}$2
Torso Gravity Error $1-\cos\theta_{\text{torso}}$ $\cos\theta_{\text{torso}}=\frac{\mathbf{g}^{\text{torso}}\cdot\mathbf{g}^{\text{target}}}{\|\mathbf{g}^{\text{torso}}\|\|\mathbf{g}^{\text{target}}\|}$2
Knee Gravity Error $\frac{1}{2}\Big{(}(1-\cos\theta_{\text{left}}^{\text{knee}})+(1-\cos\theta_{\text{right}}^{\text{knee}})\big{)}\Big{)},$ $\cos\theta^{\text{knee}}=\frac{\mathbf{g}^{\text{knee}}\cdot\mathbf{g}^{\text{target}}}{\|\mathbf{g}^{\text{knee}}\|\|\mathbf{g}^{\text{target}}\|}$2
Figure 7: Getting-up from prone pose result visualization of Tao et al. [70]. The motion generated by method [70] is highly unstable and unsafe, and it keeps jittering and jumping during the getting-up phase.
Appendix A Rewards
A.1 Rewards Components in Stage I
Detailed reward components used in Stage I are summarized in Tab. II.
A.2 Rewards Components in Stage II
Detailed reward components used in Stage II are summarized in Tab. III.